The sum of the ages of Denise and Earl is 42 years. Earl is 8 years younger than Denise. How old is each?
Found 2 solutions by sujatha.krishna, kiru_khandelwal:
Answer by sujatha.krishna(11)

Lets assume -
Denise = X and Earl = Y
X+Y=42 (Sum of their ages is 42)
-X+Y=-8 (This is actually Y+8=X which is translated from Earl is 8 years younger than Denise)
--------
2Y=34
--------
Y = 17
X = 42-17 = 25
So Earl is 17 and Denise is 25 !!

Answer by kiru_khandelwal(79)

Let Denis age be x
and Earl age be y
Sum of their ages is 42
=> x+y = 42..........equation(1)
Earl is 8 years younger then Denis
=> y = x -8............equation(2)
Substituting the value of y from equation(2) to equation(1)
=> x+y = 42
=> x + (x-8) = 42
=> 2x - 8 = 42
=> 2x = 42 +8
=> 2x = 50
=> x = 25
So Denis age is 25
Earl is 8 years younger then denis
SO Earl age = 25-8 = 17